# Assessing the Effects of Monetary Shocks on Macroeconomic Stars: A SMUC-IV Framework

### v1

@@ 4.2 Estimating the Marginal Likelihood for SMUC-IV
We now outline our conditional Monte Carlo improved modified harmonic mean estimator for the marginal likelihood of the SMUC-IV. In our modeling setting, the marginal likelihood is given by
@@
The second equality follows from the conditional independence structure of the prior distributions. In the last equality, we have integrated out the state parameters $\bm{\tau}$ and hyperparameters $(\kappa_{1},\kappa_{2})$. Specifically, we have:
@@
Note that the expression in (21) can be obtained directly from result given in (20), which shows that the joint conditional distribution of $(\mathbf{y},\bm{\tau})$ is Gaussian. Consequently, $p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)$ can be derived using standard properties of the Gaussian distribution. See Appendix B for more details.
Using the analytical expressions for (21) and (22), we can estimate the marginal likelihood of our SMUC-IV with the following conditional Monte Carlo improved modified harmonic mean estimator:
@@
where $(\bm{\Phi}^{(i)},\mathbf{B}^{(i)},\alpha^{(i)},\beta^{(i)})$, $i=1,\ldots,R$, are posterior draws that can be obtained using the posterior sampler described in the last section. Note that the key difference between this estimator and the standard modified harmonic mean estimator is the use of the likelihood function $p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)$, which is unconditional on the high-dimensional state parameters $\bm{\tau}$, and the marginal prior $p(\bm{\Phi})$. This formulation substantially reduces the dimensionality of the numerical integration, which in turn reduces the Monte Carlo variance of the estimator, resulting in greater numerical stability and estimation precision (see Chan, (2023) for more discussion).
We now turn to the choice of the tuning density function $q(\bm{\Phi},\mathbf{B},\alpha,\beta)$, a critical component of our estimator (23) that is essential for the accuracy of marginal likelihood estimation. Although the posterior density is the theoretically optimal choice, it is computationally intractable. Therefore, we follow Chan, (2023) and approximate the posterior using a truncated Gaussian density. Specifically, we consider a tuning density function that takes the following form:
@@
where each $q_{j}(\cdot)$ for $j\in\{\Phi,B,\alpha,\beta\}$ is a truncated Gaussian density with mean and covariance matrix set to the estimated posterior mean and covariance matrix. [^5] Using an appropriately truncated tuning density can ensure that the modified harmonic mean estimator has a finite variance (Geweke, , 1999). The full details for $q_{\Phi}(\bm{\Phi})$, $q_{B}(\mathbf{B})$, $q_{\alpha}(\alpha)$, and $q_{\beta}(\beta)$ are provided in Appendix B.
@@ 5 Empirical Analysis
In this section, we first describe the data in Section 5.1 and then conduct a model comparison exercise in Section 5.2 to validate our proposed SMUC-IV framework. Section 5.3 presents our estimates of the macroeconomic stars—namely, the level of potential output, the growth rate of potential output, trend inflation, and the neutral interest rate. Section 5.4 investigates the effects of monetary policy shocks on these stars. Subsequently, Section 5.5 assesses the role of the monetary policy shocks in driving the evolution of the macroeconomic stars. Finally, robustness analyses are reported in Section 5.6.
@@ 5.1 Data and External Instrument
We estimate the empirical model using the following quarterly data from 1987:Q4 to 2023:Q4: real GDP, the GDP deflator, the federal funds effective rate, the shadow interest rate, and the orthogonal monetary policy surprise (Bauer and Swanson, , 2023). To capture both conventional and unconventional monetary policy when the nominal federal funds rate is at the effective lower bound (ELB), we use the Wu and Xia, (2016) shadow short rate. Consistent with the quarterly frequency of our data we use four lags in our model. We transform real GDP as $100\times\log(x_{t})$ and compute the quarterly GDP deflator as annualized log growth, $400\times\log(x_{t}/x_{t-1})$. The federal funds effective rate, shadow interest rate, and orthogonal monetary policy surprise are originally monthly series; we obtain their quarterly versions by averaging the three monthly observations within each quarter. Real GDP, the GDP deflator, and the federal funds effective rate are available from the FRED database. We consider different measures of inflation and the interest rate as well as consider the bond premium (Favara et al., , 2016) as a control variable in the vector autoregression, see section 5.6. The shadow interest rate is from the website of the Federal Reserve Bank of Atlanta, the excess bond premium is from the website of the Board of Governors of the Federal Reserve System, and the orthogonal monetary policy surprise series is from the website of the Federal Reserve Bank of San Francisco.
In this paper, we use the orthogonal monetary policy surprise series developed by Bauer and Swanson, (2023) as our external instrument. This series identifies monetary policy shocks from high-frequency asset price movements in narrow windows around policy announcements. High-frequency interest rate changes around FOMC announcements are a common tool for identifying monetary policy effects, but recent studies have questioned their exogeneity and relevance as instruments, particularly for estimating macroeconomic impacts (e.g., Ramey, , 2016; Miranda-Agrippino and Ricco, , 2021). For instance, monetary policy surprises may be correlated with publicly available macroeconomic and financial data released before FOMC announcements. To address these concerns, Bauer and Swanson, (2023) (i) expand the set of monetary policy events to include speeches by the Federal Reserve Chair—roughly doubling the number of announcements—and (ii) orthogonalize the resulting surprises with respect to pre-announcement macroeconomic and financial data to account for predictability via the “Fed response to news” channel.
@@ 5.2 Model Comparison Exercise
The first alternative specification imposes the restriction that monetary policy shocks have no contemporaneous effects on all of the trend components, that is, the macroeconomic stars of interest in this paper. We refer to this model as SMUC-IV-R1. Next, we consider a restricted version of our SMUC-IV that imposes zero correlation between the external instrument and all structural shocks by setting $\beta=0$. A comparison between our proposed SMUC-IV and SMUC-IV-R2 serves to test the relevance condition of the instrument. Note that under the restriction $\beta=0$, the proxy equation (11) is independent of the system of equations in the unobserved components model specified in (9). Therefore, by imposing various patterns of zero restrictions on $\mathbf{B}$, we can assess different correlation structures between the trend and cycle innovations. In our model comparison exercise, we further consider two nested versions of SMUC-IV-R2. The first nested version of SMUC-IV-R2 is denoted as SMUC-IV-R3. This specification assumes that innovations between the trend and cycle components are uncorrelated, while allowing for correlation within the innovations of the trend components and within the innovations of the cycle components, respectively. Specifically, under SMUC-IV-R3 we assume the $6\times 6$ contemporaneous response matrix $\mathbf{B}$ to be a block diagonal matrix with each block of dimension $3\times 3$. For the second nested version of SMUC-IV-R2, we assume $\mathbf{B}$ to be a diagonal matrix and denote this model as SMUC-IV-R4. This specification is similar to that in many conventional studies on multivariate unobserved components models, which assumes all innovations of the trend and cycle components are mutually independent. Table 1 provides a summary of the competing models in our model comparison exercise.
@@
Table 2 reports the log marginal likelihood estimates. The results provide significant evidence that our proposed SMUC-IV model outperforms all alternative specifications considered. A few findings are also worth highlighting. First, by comparing the SMUC-IV with the SMUC-IV-R1, the difference in the log marginal likelihoods is about $13$, strongly supporting the non-zero contemporaneous effects of monetary policy shocks on the trend components. Second, the SMUC-IV-R2 performs better than the SMUC-IV-R3 and SMUC-IV-R4. This suggests that allowing for correlated innovations among all trend and cycle components is an important modeling feature supported by the data. In particular, the log marginal likelihood of SMUC-IV-R2 is about $24$ higher than that of SMUC-IV-R3, indicating that the assumption of independent innovations for the trend and cycle components is empirically implausible in our empirical analysis. Lastly, the SMUC-IV outperforms SMUC-IV-R1, suggesting that the relevance condition of the instrument is satisfied, although the difference in the log marginal likelihoods between these two models is only about 33.666In Section 5.6, we also conduct a robustness analysis by considering an alternative instrument in our application, and we find that our main empirical results remain robust under this alternative choice of instrument.
@@ 5.3 Estimates of Macroeconomic Stars
Although macroeconomic stars—for example, the neutral interest rate—have been defined in various ways across the literature, most theoretical frameworks suggest that shifts in the low-frequency components of macroeconomic variables are closely linked to movements in any theoretically defined star. Accordingly, we examine the persistent evolution of macroeconomic stars through the lens of trend dynamics. Using US data, we jointly estimate four stars: the level of potential GDP, $g^{*}$, the growth rate of potential GDP, $\Delta g^{*}$, and trend inflation, $\pi^{*}$, and the neutral interest rate, $r^{*}$. These stars are measured as long-run trends in the spirit of the Beveridge-Nelson decomposition (Beveridge and Nelson, , 1981).
@@
Panel (a) of Figure 1 presents the estimates of $^{*}. gg^{*}$ rises over time but experiences two notable drops during the 2007–2009 Global Financial Crisis and the COVID-19 pandemic. In both cases, the level of $g^{*}$ fails to return to its pre-crisis trajectory, suggesting that transitory shocks may translate into permanent reductions in the long-run growth path.
Panel (b) of Figure 1 shows estimates of $\Delta g^{*}$. From the early to late 1990s, $\Delta g^{*}$ rises noticeably, likely reflecting the internet technology boom. After the 1990s, however, it follows a downward trend until around 2010. Then it follows an upward trend until the early 2020s. Our estimates show a similar tendency to those in (Grant and Chan, 2017a, ; Grant and Chan, 2017b, ; Maffei-Faccioli, , 2025), albeit with considerable fluctuations. Sharp declines are evident during major crises such as the dot-com bust in the early 2000s, the Global Financial Crisis of 2007–2009, and the COVID-19 pandemic. These episodes coincide with well-documented hysteresis effects, where severe recessions leave lasting scars on the economy’s productive capacity (Cerra and Saxena, , 2005, 2008; Cerra et al., , 2023).
Panel (c) of Figure 1 reports the estimates of $\pi^{*}$. Consistent with prior work (e.g., Stock and Watson, , 2007, 2016; Chan et al., , 2018; Eo et al., , 2023), $\pi^{*}$ displays remarkable stability: after falling in the 1990s, it remains anchored near 2% from the 2000s until the pandemic period. This pattern highlights the Federal Reserve’s success in stabilizing long-run inflation expectations.
Finally, Panel (d) of Figure 1 presents estimates of $r^{*}$. The estimates show a steady decline from the 1990s onward, followed by relative stability during the 2010s-2020s. Our estimates of $r^{*}$ capture the secular decline in the neutral interest rate documented in the literature on estimating $r^{*}$ (e.g., Laubach and Williams, , 2003; Lubik and Matthes, , 2015; Del Negro et al., , 2017; Morley et al., , 2024).
@@
5.4 Do Monetary Policy Shocks Affect Macroeconomic Stars?
Having obtained plausible estimates of the macroeconomic stars, we now turn to the paper’s central aim: exploring the effects of monetary policy shocks on these stars. Specifically, we examine how monetary policy influences $\Delta g^{*}$, $\pi^{*}$, and $r^{*}$. Monetary policy shocks are identified using an external instrument based on high-frequency policy surprises developed by Bauer and Swanson, (2023).
@@ 5.4 Do Monetary Policy Shocks Affect Macroeconomic Stars?
Figure 2 shows the posterior distributions of the impulse responses of $\Delta g^{*}$, $\pi^{*}$, and $r^{*}$ to a one-standard-deviation contractionary monetary policy shock. This shock raises the median response of the interest rate by about 0.29 basis points on impact (see Figure 3). By construction, the responses are constant over time, as the model assumes the stars follow a random walk.
Panel (a) of Figure 2 shows that $\Delta g^{*}$ falls by about 0.1 percentage point. The negative effects on the growth rate of potential output align with the main finding of Jordà et al., (2024) that monetary policy can have long-run effects on the real economy, suggesting an innovation channel whereby monetary tightening curbs spendings on R&D and innovation, limits productivity gains, and reduces productive investment, thereby depressing the growth rate of potential output (see, e.g., Stadler, , 1990; Moran and Queralto, , 2018; Ma and Zimmermann, , 2023; Fornaro and Wolf, , 2023; Meier and Reinelt, , 2024).
Panel (b) of Figure 2 shows that contractionary monetary policy shocks also reduce $\pi^{*}$ by about 0.2 percentage points. At first glance, significant effects of monetary policy shocks on trend inflation may appear undesirable, as they seem to imply de-anchoring of long-run inflation expectations. However, when long-run expectations are persistently off target, such effects may reflect a re-anchoring channel, in which policy deliberately shifts expectations back toward the target. As shown by our counterfactual analysis (see Figure 4), monetary policy shocks pull trend inflation down make it move around the 2% target from mid-1990s to the onset of Covid-19 pandemic. This finding is consistent with evidence that such shocks lower survey-based long-run inflation expectations, pointing to a re-anchoring channel in which monetary policy shifts long-run inflation expectations back toward target when they drift persistently away from it (see, e.g., Jarociński and Karadi, , 2020; Diegel and Nautz, , 2021).
Panel (c) shows that $r^{*}$ falls by about 0.3 basis points. This decline, caused by contractionary monetary policy shocks, may operate through two complementary mechanisms. First, the innovation channel in panel (a) indicates that contractionary monetary policy shocks lowers potential output growth; because trend growth is a key determinant of the neutral rate (e.g., Laubach and Williams, , 2003), this directly pushes $r^{*}$ down. Second, the re-anchoring channel in panel (b) indicates that these shocks lower $\pi^{*}$, thereby raising the real interest rate and also encouraging firms to cut spending on R&D and productive investment —reinforcing the innovation channel and further depressing $r^{*}$.
A more practical question concerns how actual macroeconomic variables respond to contractionary monetary policy shocks once policy-induced shifts in the macroeconomic stars are taken into account. Figure 3 shows that the impulse responses of the actual variables align with standard macroeconomic predictions: the nominal interest rate rises, output falls, and inflation declines. Panel (a) of Figure 3 shows that output falls by about 1.5 percent after five years. Panel (b) indicates that inflation declines by roughly 0.17 percentage points after five years, following the 0.29-basis-point increase in the nominal interest rate shown on impact in Panel (c). Notably, when monetary policy shocks are allowed to affect the macroeconomic stars, the responses of the actual variables become markedly more persistent.
@@
Figure 3: The plots shows the impulse responses with 68% impulse responses of the actual variable to a monetary policy shock.
5.5 Are the Monetary Policy Shocks a Key Driver for Shaping the Macroeconomic Stars?

### v2

@@ 4.2 Estimating the Marginal Likelihood for SMUC-IV
We now outline our conditional Monte Carlo improved modified harmonic mean estimator for the marginal likelihood of the SMUC-IV. In our modelling setting, the marginal likelihood is given by
@@
The second equality follows from the conditional independence structure of the prior distributions. In the last equality, we have integrated out the state parameters $\bm{\tau}$ and the hyperparameters $(\kappa_{1},\kappa_{2})$. Specifically, we have:
@@
Note that the expression in (21) can be obtained directly from the result in (20) by using the property that the marginal distribution of joint Gaussian variates is Gaussian. To be specific, let $\bm{\Lambda}=\mathbf{K}^{-1}$ be the covariance matrix of the joint conditional distribution of $(\bm{\tau}^{\prime},\mathbf{y}^{\prime})^{\prime}$, and partition $\bm{\Lambda}$ according to the dimensions of $\bm{\tau}$ and $\mathbf{y}$ as
A standard property of the Gaussian distribution implies that the marginal distribution of $\mathbf{y}$ (of dimension $4T\times 1$) is a Gaussian distribution with mean vector $\bm{\mu}_{y}$ and covariance matrix $\bm{\Lambda}_{y}$. This yields
For the analytical expression of the marginal prior $p(\bm{\Phi})$ in (22), we refer readers to Appendix B for details.
Given the analytical expressions for (21) and (22), we can estimate the marginal likelihood of our SMUC-IV with the following conditional Monte Carlo improved modified harmonic mean estimator:
@@
where $(\bm{\Phi}^{(i)},\mathbf{B}^{(i)},\alpha^{(i)},\beta^{(i)})$, $i=1,\ldots,R$, are posterior draws that can be obtained using the posterior sampler described in the last section. Note that the key difference between this estimator and the standard modified harmonic mean estimator is the use of the likelihood function $p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)$, which is unconditional on the high-dimensional state parameters $\bm{\tau}$, and the marginal prior $p(\bm{\Phi})$. This formulation substantially reduces the dimensionality of the numerical integration, which in turn reduces the Monte Carlo variance of the estimator, resulting in greater numerical stability and estimation precision (see chan2023comparing for more discussion).
We now turn to the choice of the tuning density function $q(\bm{\Phi},\mathbf{B},\alpha,\beta)$, a critical component of our estimator (23) that is essential for the accuracy of marginal likelihood estimation. Although the posterior density is the theoretically optimal choice, it is computationally intractable. Therefore, we follow chan2023comparing and approximate the posterior using a truncated Gaussian density. Specifically, we consider a tuning density function that takes the following form:
@@
where each $q_{j}(\cdot)$ for $j\in\{\Phi,B,\alpha,\beta\}$ is a truncated Gaussian density with mean and covariance matrix set to the estimated posterior mean and covariance matrix. [^6] Using an appropriately truncated tuning density can ensure that the modified harmonic mean estimator has a finite variance (geweke1999using). The full details for $q_{\Phi}(\bm{\Phi})$, $q_{B}(\mathbf{B})$, $q_{\alpha}(\alpha)$, and $q_{\beta}(\beta)$ are provided in Appendix B.
@@ 5 Empirical Analysis
In this section, we first describe the data in Section 5.1 and then conduct a model comparison exercise in Section 5.2 to validate our proposed SMUC-IV framework. In particular, this exercise provides strong evidence in favor of the influence of monetary policy shocks on the stars. Section 5.3 presents our estimates of the macroeconomic stars—namely, the potential GDP growth, trend inflation, and the neutral interest rate. Section 5.4 examines the effects of monetary policy shocks on these stars. Subsequently, Section 5.5 assesses the role of the monetary policy in driving the evolution of the macroeconomic stars. Section 5.6 provides a discussion of the results and their policy implications. Finally, robustness analyses are reported in Section 5.7.
@@ 5.1 Data and External Instrument
We estimate the empirical model using the following quarterly data from 1988:Q1 to 2023:Q3: real GDP, the GDP deflator, the federal funds effective rate, the shadow interest rate, and the orthogonal monetary policy surprise (bauer2023reassessment). To capture both conventional and unconventional monetary policy when the nominal federal funds rate is at the effective lower bound (ELB), we use the wu2016measuring shadow short rate. Consistent with the quarterly frequency of our data, we use four lags in our model. We transform real GDP as $100\log(x_{t})$ and compute real GDP growth as the first difference of this series. The quarterly GDP deflator inflation rate is measured as annualized log growth, 400 $\log(x_{t}/x_{t-1})$. The federal funds effective rate, shadow interest rate, and orthogonal monetary policy surprise are originally monthly series; we obtain their quarterly versions by averaging the three monthly observations within each quarter. Real GDP, the GDP deflator, and the federal funds effective rate are available from the FRED database. We consider different measures of inflation and the interest rate as well as consider the bond premium (favara2016updating) as a control variable in the vector autoregression, see section 5.7. The shadow interest rate is from the website of the Federal Reserve Bank of Atlanta, the excess bond premium is from the website of the Board of Governors of the Federal Reserve System, and the orthogonal monetary policy surprise series is from the website of the Federal Reserve Bank of San Francisco.
In this paper, we use the orthogonal monetary policy surprise series developed by bauer2023reassessment as our external instrument. This series identifies monetary policy shocks from high-frequency asset price movements in narrow windows around policy announcements. High-frequency interest rate changes around FOMC announcements are a common tool for identifying monetary policy effects, but recent studies have questioned their exogeneity and relevance as instruments, particularly for estimating macroeconomic impacts (e.g., ramey2016macroeconomic; miranda2021transmission). For instance, monetary policy surprises may be correlated with publicly available macroeconomic and financial data released before FOMC announcements. To address these concerns, bauer2023reassessment (i) expand the set of monetary policy events to include speeches by the Federal Reserve Chair—roughly doubling the number of announcements—and (ii) orthogonalize the resulting surprises with respect to pre-announcement macroeconomic and financial data to account for predictability via the “Fed response to news” channel. In Section 5.7, we also conduct a robustness analysis by considering an alternative instrument in our application, and we find that our main empirical results remain robust under this alternative choice of instrument.
@@ 5.2 Model Comparison Exercise
The first alternative specification imposes the restriction that monetary policy shocks have no contemporaneous effects on all of the trend components, that is, the macroeconomic stars of interest in this paper. We refer to this model as SMUC-IV-R1. Next, we consider a restricted version of our SMUC-IV model that imposes zero correlation between the external instrument and all structural shocks by setting $\beta=0$; we refer to this specification as SMUC-IV-R2. A comparison between our proposed SMUC-IV and SMUC-IV-R2 serves to test the relevance condition of the instrument. Note that under the restriction $\beta=0$, the proxy equation (11) is independent of the system of equations in the unobserved components model specified in (9). Therefore, by imposing various patterns of zero restrictions on $\mathbf{B}$, we can assess different correlation structures between the trend and cycle innovations. In our model comparison exercise, we further consider two nested versions of SMUC-IV-R2. The first nested version of SMUC-IV-R2 is denoted as SMUC-IV-R3. This specification assumes that innovations between the trend and cycle components are uncorrelated, while allowing for correlation within the innovations of the trend components and within the innovations of the cycle components, respectively. Specifically, under SMUC-IV-R3 we assume the $6\times 6$ contemporaneous response matrix $\mathbf{B}$ to be a block diagonal matrix with each block of dimension $3\times 3$. For the second nested version of SMUC-IV-R2, we assume $\mathbf{B}$ to be a diagonal matrix and denote this model as SMUC-IV-R4. This specification is similar to that in many conventional studies on multivariate unobserved components models, which assumes all innovations of the trend and cycle components are mutually independent. Table 1 provides a summary of the competing models in our model comparison exercise.
@@
Table 2 reports the log marginal likelihood estimates. The results provide significant evidence that our proposed SMUC-IV model outperforms all alternative specifications considered. A few findings are also worth highlighting. First, by comparing the SMUC-IV with the SMUC-IV-R1, the difference in the log marginal likelihoods is about $18$, strongly supporting the non-zero contemporaneous effects of monetary policy shocks on the trend components. Second, the SMUC-IV-R2 performs better than the SMUC-IV-R3 and SMUC-IV-R4. This suggests that allowing for correlated innovations among all trend and cycle components is an important modelling feature supported by the data. In particular, the log marginal likelihood of SMUC-IV-R2 is about $22$ higher than that of SMUC-IV-R4, indicating that the assumption of independent innovations for the trend and cycle components is empirically implausible in our empirical analysis. Lastly, the SMUC-IV outperforms SMUC-IV-R2, suggesting that the relevance condition of the instrument is satisfied.
@@ 5.3 Estimates of Macroeconomic Stars
Although macroeconomic stars—for example, the neutral interest rate—have been defined in various ways across the literature, most theoretical frameworks suggest that shifts in the low-frequency components of macroeconomic variables are closely linked to movements in any theoretically defined star. Accordingly, we examine the persistent evolution of macroeconomic stars through the lens of trend dynamics. Using US data, we jointly estimate three stars: the potential GDP growth, $g^{*}$, the trend inflation, $\pi^{*}$, and the neutral interest rate, $r^{*}$.
@@
Panel (a) of Figure 2 shows estimates of $g^{*}$. From the early to late 1990s, $g^{*}$ rises noticeably, likely reflecting the internet technology boom. After the 1990s, however, it follows a downward trend until around 2010. Then it follows an upward trend until the early 2020s. Our estimates show a similar tendency to those in (grant2017bayesian; grant2017reconciling; maffei2025identifying). Sharp declines are evident during major crises such as the dot-com bust in the early 2000s, the Global Financial Crisis of 2007–2009, and the COVID-19 pandemic. These episodes coincide with well-documented hysteresis effects, where severe recessions leave lasting scars on the economy’s productive capacity (cerra2005did; cerra2008growth; cerra2023hysteresis).
Panel (b) of Figure 2 reports the estimates of $\pi^{*}$. Consistent with prior work (e.g., stockwatson2; stock2016core; chan2018new; eo2023understanding), $\pi^{*}$ displays remarkable stability: after falling in the 1990s, it remains anchored near 2% from the 2000s until the pandemic period. This pattern highlights the Federal Reserve’s success in stabilizing long-run inflation expectations.
Finally, Panel (c) of Figure 2 presents estimates of $r^{*}$. The estimates show a steady decline from the 1990s onward, followed by relative stability during the 2010s-2020s. Our estimates of $r^{*}$ capture the secular decline in the neutral interest rate documented in the literature on estimating $r^{*}$ (e.g., laubach2003measuring; lubik2015calculating; del2017safety; morley2024simple).
We also report the estimated cyclical components in Figure 3. Overall, the cycles appear stationary, suggesting that our proposed model effectively decomposes the time series of interest into trends and cycles. The cyclical components capture major economic fluctuations, including the recessions in the early 1990s, the global financial crisis of 2008-2009, and the sharp contraction during the COVID-19 pandemic in 2020.
Figure 3: The plot shows cycle estimates with 68% credible bands.
5.4 The Effects of Monetary Policy Shocks on the Macroeconomic Stars
Having obtained plausible estimates of the macroeconomic stars, we now examine the effects of monetary policy shocks on $g^{*}$, $\pi^{*}$, and $r^{*}$. We identify monetary policy shocks using an external instrument based on the high-frequency policy surprises constructed by bauer2023reassessment.
@@ 5.4 The Effects of Monetary Policy Shocks on the Macroeconomic Stars
Figure 4 shows the posterior distributions of the impulse responses of $g^{*}$, $\pi^{*}$, and $r^{*}$ to a one-standard-deviation contractionary monetary policy shock. This shock raises the median response of the nominal interest rate by about 29 basis points on impact (see Figure 5(f)). By construction, the responses are constant over time, as the model assumes the stars follow a random walk.
Panel (a) of Figure 4 shows that $g^{*}$ falls by about 0.05 percentage points. The negative effects on potential GDP growth align with the empirical evidence in moran2018innovation, garga2021output, ma2023monetary, jorda2024long, and meier2024monetary, which suggests that tighter monetary policy restrains innovation investment and, in turn, productivity. Panel (b) of Figure 4 shows that contractionary monetary policy shocks also reduce $\pi^{*}$ by about 0.2 percentage points. The decline in trend inflation is consistent with evidence that contractionary monetary policy shocks reduce long-term inflation expectations (e.g., jarocinski2020deconstructing; diegel2021long). Panel (c) shows that $r^{*}$ falls by about 25 basis points. The decline in the neutral interest rate after contractionary monetary policy shocks could be interpreted as reflecting the adverse effects of monetary tightening on innovation investment.
As a complement to our analysis of the effects of monetary policy shocks on the stars, we also report the cyclical responses and the responses of the observed variables to contractionary monetary policy shocks in Figure 5.
@@
Figure 5: The plots show the impulse responses of the cyclical components and the observed variables to a monetary policy shock, along with 68% credible intervals.
Figure 5 shows that the impulse responses align with standard macroeconomic predictions: the nominal interest rate rises, real GDP growth falls, and inflation declines. The figure also indicates that contractionary monetary policy shocks have more persistent effects on the business cycle than expected, taking about 3–5 years to fade out. Moreover, the responses of the observed variables appear to be largely permanent.
Overall, our results suggest that monetary tightening may generate medium-run adverse effects on aggregate demand, which in turn suppress investment in innovation and weaken long-run productivity, leaving potentially permanent scars on the economy.
5.5 The Role of Monetary Policy Shocks in Shaping Macroeconomic Stars